OpenStudy (anonymous):

can someone help me??? find functionsf and g so that f composite g =H... H(x)=1/(x^2-8)

OpenStudy (anonymous):

\[g(x) = x^{2}-8, f(x) = \frac{1}{x}\] thats one way to do it.

OpenStudy (anonymous):

or: \[g(x) = x^{2}, f(x) = \frac{1}{x-8}\] that works too.

OpenStudy (anonymous):

or if you wanna be cheap about it (this will work in every case): \[g(x) = x, f(x) = \frac{1}{x^{2}-8}\]

OpenStudy (anonymous):

how do u got that??? can u show me step by step???

OpenStudy (anonymous):

sure, although its not something i can really show, its a bunch of explanation. So this is your problem: \[\frac{1}{x^{2}-8}\] First, i want to see what is the general form of this function. Its general form is: \[\frac{1}{junk}\] That is why i mae the outside function (f in your case): \[f(x) = \frac{1}{x}\]

OpenStudy (anonymous):

Second, i ask myself, "what is the junk?", well we can see that the junk is: \[x^{2}-8\] so thats what i made the inside function. g has to be: \[g(x) = x^2-8\]

OpenStudy (anonymous):

So following these two steps, could you tell me what two functions would form: \[\frac{1}{x+5}\] ? its the exact same process.

OpenStudy (anonymous):

1/x and x+5.. is that right?

OpenStudy (anonymous):

yep, perfect :)

OpenStudy (anonymous):

sometimes it might get a little more complicated. it just comes down to can you identify the junk.

OpenStudy (anonymous):

ok... thnkas for helping..

OpenStudy (anonymous):

hiihih ok

OpenStudy (anonymous):

a slightly harder example: \[(x+3)^{2} +5(x+3) -2\] this is in the form: \[(junk)^2+5(junk)-2\] so my two functions would be: \[f(x) = x^2+5x-2, g(x) = x+3\]

OpenStudy (anonymous):

good job, medals for everyone lol